I have no Idea were to start with this problem, can you please help with this function?
In 1990, the life expectancy of males in a certain country was 70.2 years. In 1997, it was 72.7 years. Let E represent the life expectancy in year t and let t represent the number of years since 1990.
The linear function E(t) that fits the data is
so fair I came up with
E(t)= _t +_.
please help!!

Answer by stanbon(75887)

In 1990, the life expectancy of males in a certain country was 70.2 years. In 1997, it was 72.7 years. Let E represent the life expectancy in year t and let t represent the number of years since 1990.
The linear function E(t) that fits the data is
-----------------------------------------------
You have two points relating (t,E(t)): (90,70.2) and (97,72.7)
slope = (72.7-70.2)/(7-0) = 2.5/7 = 5/14
----------------
Equation form is y = mx+b; you have x,y,and m; solve for "b" as follows:
70.2 = (5/14)*0 + b
b = 70.2
-------------------------
EQUATION:
E(t)= (5/14)t + 70.2
=================================
Cheers,
Stan H.